OpenStudy (anonymous):

Solve absolute value 2|3x + 1| − 1 = 29. Thanks in advance

OpenStudy (zzr0ck3r):

first add 1 to both sides, and then divide by 2 to isolate the |3x+1| term.

OpenStudy (zzr0ck3r):

tell me what you have

OpenStudy (anonymous):

I got 4.6 and -5

OpenStudy (zzr0ck3r):

um try what I told you

OpenStudy (zzr0ck3r):

2|3x + 1| − 1 = 29 add 1 to both sides 2|3x+1|=30 divide by 2 |3x+1|=15

OpenStudy (zzr0ck3r):

do you see?

OpenStudy (anonymous):

yeah I got that then I solve right?

OpenStudy (zzr0ck3r):

ok to solve |a|=b ---> a=-b and a=b

OpenStudy (zzr0ck3r):

so you have 3x+1=15 and 3x+1=-15

OpenStudy (zzr0ck3r):

is this what you did?

OpenStudy (anonymous):

yeah 3x+1=15 =4.66

OpenStudy (anonymous):

the other one though I did wrong for some reason

OpenStudy (anonymous):

i had 3x+1=-14

OpenStudy (zzr0ck3r):

done write it like this 3x+1=15 =4.66 that implies 15=4.66

OpenStudy (zzr0ck3r):

but I get what your saying

OpenStudy (zzr0ck3r):

ok you got it from here then?

OpenStudy (zzr0ck3r):

the other is -16/3

OpenStudy (anonymous):

-5.333 yes

OpenStudy (anonymous):

you see what i did was solve from up top as 2 I3x+1I -1=29 and 2 I3x+1I -1=-29

OpenStudy (anonymous):

meh stupid software would've counted it as wrong still. thanks!

OpenStudy (zzr0ck3r):

don't use decimal, write -16/3